# Fair Graph Machine Learning under Adversarial Missingness Processes

### v4

@@ 3 Introducing Bias via an Adversarial Missingness Process
The AMMBD problem is NP-hard.
See proof of Theorem 1 in Appendix A. The idea of our proof is to frame AMMBD as an adversarial version of active learning where the goal is to strategically minimize the accuracy of the model for certain nodes by selecting labels. More specifically, we apply a popular formulation of active learning as a coverage problem yehuda2022active ; ren2021survey . The adversary will target nodes either (1) with sensitive attribute $s_{v}=1$ and in the undesired class $y_{v}=0$ or (2) with sensitive attribute $s_{v}=1$ and in the desired class $y_{v}=1$. Targeting these nodes will decrease the bias in the imputed data.
Theorem 1 can be interpreted as a positive result, as it shows that, in theory, a simpler surrogate of the adversary’s objective is hard to optimize. The best-known approximation for the minimum k-union problem chlamtavc2017minimizing used in our reduction is $O(n^{1/4+\epsilon})$, where $n$ is the number of sets, which is mapped to the size of the graph in our problem. This means that approximately $k/n^{1/4}$ nodes would have to be selected to achieve the optimal coverage for $k$ nodes. However, we show empirically that a simple heuristic based on node degrees often achieves good results for the problem when an independent imputation model $f_{I}$ based on a GNN is applied.
Heuristic for AMMBD: We generate missing values synthetically in our experiments. To simulate an adversarial missingness process, we select low-degree nodes to have missing values. This choice is based on previous work showing that low-degree nodes have lower prediction accuracy tang2020investigating ; liu2023generalized ; ju2024graphpatcher . To generate $k$ nodes with missing sensitive data, we choose the $k$ nodes with the lowest degree.
To see the impact of adversarial missingness, we compare the imputation of adversarial missingness and random missingness. We show the correlation of sensitive information and true class labels for different datasets in Figure 2. We use a Graph Convolutional Network (GCN) to predict the missing sensitive values for independent imputation. Figures 2(a) and 2(b) show the correlation with adversarial missingness and random missingness, respectively. When $50\%$ sensitive information is available, the predicted value with independent imputation underrepresents the true underlying bias in all datasets with adversarial missingness. For random missingness, the performance is slightly better. Figure 2 (c) and (d) show results for BFtS, which will be introduced in the next section. BFtS rarely underestimates the true bias in the datasets. Our results provide strong evidence that off-the-shelf imputation methods often misguide fair graph machine learning algorithms due to underestimated bias, especially with adversarial missingness processes.
Figure 2: Empirical results showing that a simple degree-based adversarial missingness process is effective at minimizing the bias of an independent imputation model (GNN) compared with a random missingness process. The last column of each matrix shows the true bias in the data, and lower values show that the bias is underestimated. We also show results for our approach (BFtS) described in Section 4 that addresses this problem using a 3-player adversarial imputation method.
As our proposed model operates under a worst-case bias assumption, it may overestimate the bias in the complete data, as illustrated in Figure 2 using the Bail dataset. This results in a trade-off between fairness and utility, which is governed by a hyperparameter $\beta$. Figures 3(a), (b), and (c) show the F1 score, $1-\Delta\text{DP}$, and $1-\Delta\text{EQOP}$, respectively, as $\beta$ varies and with $30\%$ of sensitive values observed. Here, F1 serves as the utility metric, while $1-\Delta\text{DP}$ and $1-\Delta\text{EQOP}$ quantify fairness (see Section 5). We compare the performance of our BFtS model against a fair adversarial model trained with complete data. Adjusting $\beta$ enables controlling the balance between utility and robustness to worst-case bias, with appropriate tuning yielding an effective trade-off.
4 Fairness-Aware Adversarial Missing Data Imputation for Graphs
@@
We introduce a framework called BFtS (Better Fair than Sorry) that allows fair GNN training with adversarial missing sensitive data. BFtS applies a 3-player adversarial learning scheme, where the fair GNN is trained jointly with two adversaries—one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness. As a result, fairness assessments by the GNN are based on the worst-case imputed values from the adversarial imputation.
@@ 4.1 Proposed Model (BFtS)
Figure 4: 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{C}$ generates node representations by minimizing the classification loss $\mathcal{L}_{C}$ (Eqn. 1) and the maximizing sensitive attribute prediction loss $\mathcal{L}_{A}$ (Eqn. 3). $f_{A}$ predicts sensitive attributes using representations from $f_{C}$ by minimizing $\mathcal{L}_{A}$. $f_{I}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{I}$ (Eqn. 2) and maximizing $\mathcal{L}_{A}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are predictions from $f_{C}$, $f_{I}$ and $f_{A}$, respectively.
Our goal is to learn a fair and accurate GNN $\hat{y}=f_{C}(\mathcal{G},\mathcal{X})$ for node classification with missing sensitive values. The proposed solution is model-agnostic and based on adversarial learning. It formulates the imputation model as a second adversary of the GNN. Figure 4 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{I}$, a node classification GNN $f_{C}$, and a sensitive attribute Deep Neural Net (DNN) prediction $f_{A}$. $f_{C}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{A}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{C}$ to assess the bias. More specifically, it $f_{C}$ is assumed to be biased if the adversary $f_{A}$ can accurately predict the sensitive attribute information from the representations of $f_{C}$. To train $f_{C}$, we need to know the values of the missing sensitive attributes. The model $f_{I}$ predicts the missing sensitive attributes by taking $\mathcal{X}$ and $\mathcal{G}$ as inputs and generating the missing sensitive attributes $\hat{si}$. The goal of $f_{I}$ is to generate sensitive values that minimize the fairness of $f_{C}$ and, thus, it works as a second adversary to $f_{C}$.
@@ 4.1.1 Player architectures
GNN classifier $f_{C}$:
Node classification model $\hat{y}_{v}=f_{C}(x_{v},\mathcal{G})$ implemented using a GNN. We assume that $f_{C}$ does not apply sensitive attributes but uses other attributes in $\mathcal{X}$ that might be correlated with sensitive ones. The goal of $f_{C}$ is to achieve both accuracy and fairness. To improve fairness, $f_{C}$ tries to minimize the loss of adversary $f_{A}$.
Sensitive attribute predictor (Adversary 1) $f_{A}$:
Neural network that applies representations $\mathbf{h}_{v}$ generated by $f_{C}$ to predict sensitive attributes as $\hat{sa}_{v}=f_{A}(\mathbf{h}_{v})$. $f_{C}$ is fair if $f_{A}$ performs poorly.
Missing data imputation GNN (Adversary 2) $f_{I}$:
Predicts missing sensitive attributes as $\hat{si}_{v}=f_{I}(x_{v},\mathcal{G})$. However, besides being accurate, $f_{I}$ plays the role of an adversary to $f_{C}$ by predicting values that maximize the accuracy of $f_{A}$.
@@ 4.1.2 Loss Functions
Node Classification Loss
We apply the cross-entropy loss to learn $f_{C}$ as follows:
$\mathcal{L}_{C}=-\frac{1}{|\mathcal{V_{L}}|}\sum_{v\in\mathcal{V_{L}}}\hat{y}_{v}log(\hat{y}_{v})+(1-\hat{y}_{v})log(1-\hat{y}_{v}).$ (1)
Sensitive Attribute Imputation Loss
Because sensitive attributes tend to be imbalanced, $f_{I}$ applies the Label-Distribution-Aware Margin (LDAM) loss ldam . Let $\hat{si}=f_{I}(X,G)$ be the one-hot encoded predictions and $\hat{si}^{s}_{v}$ be the output of node $v$ for $s_{v}$. The LDAM loss of $f_{I}$ is defined as:
$\mathcal{L}_{I}=\frac{1}{|\mathcal{V_{S}}|}\sum_{v=1}^{|\mathcal{V}_{S}|}-log\frac{e^{\hat{si}^{s}_{v}-\Delta^{s}}}{e^{\hat{si}^{s}_{v}-\Delta^{s}}+\sum_{k\neq s}e^{\hat{si}^{k}_{v}}}.$ (2)
@@
The LDAM loss is a weighted version of the negative log-likelihood loss. Intuitively, since $\Delta^{j}$ is larger for smaller values of $n_{j}$, it ensures a higher margin for the smaller classes.
BFtS with no sensitive information: BFtS can generate fair outputs when very little or no sensitive information is provided by letting $f_{I}$ impute the sensitive values from the ground truth training labels $y$ using the LDAM loss. We replace $\mathcal{V}_{S}$ with $\mathcal{V}_{L}$ and $s$ with $y$ for each node $v\in\mathcal{V}_{L}$ in Eqn 2. The reasoning is that the worst-case fairness model generally assigns a more desired outcome to the non-sensitive group and a less desired outcome to the sensitive one. Therefore, the nodes predicted as the minority class will fall into the sensitive group, and vice versa. This is consistent with the worst-case assumption of BFtS. RNF du2021fairness applies a similar solution with a different loss function.
Sensitive Attribute Prediction Loss
Given the sensitive information for $\mathcal{V}_{s}$, we first replace $\hat{si}_{v}$ by ${s}_{v}$ for $v\in\mathcal{V}_{s}$ and thereby generate $\hat{s}$. The parameters of $f_{A}$ are learned using:
$\mathcal{L}_{A}=\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))]$ (3)
@@ 4.1.3 Learning the parameters of BFtS
Let $\theta_{C},\theta_{A}$, and $\theta_{I}$ be the parameters of $f_{C}$, $f_{A}$, and $f_{I}$, respectively. We learn these parameters via a 3-player adversarial learning scheme described in Figure 4. Parameters $\theta_{C}$ are optimized as follows:
$\theta_{C}^{*}=\arg\min_{\theta_{C}}\mathcal{L}_{C}+\alpha\mathcal{L}_{A}.$ (4)
@@
The parameters of the sensitive attribute predictor $\theta_{A}$ are learned by maximizing $\mathcal{L}_{A}$:
$\theta_{A}^{*}=\arg\max_{\theta_{A}}\mathcal{L}_{A}.$ (5)
To learn the parameters $\theta_{I}$ of the imputation model to generate predictions that are accurate and represent the worst-case scenario for fairness, we apply the following:
$\theta_{I}^{*}=\arg\min_{\theta_{I}}\mathcal{L}_{I}-\beta\mathcal{L}_{A}.$ (6)
@@
$\begin{split}\min_{\theta_{C}}\max_{\theta_{I},\theta_{A}}\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{A}(\mathbf{h})]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{A}(\mathbf{h}))].\end{split}$ (7)
The training algorithm and time complexity analysis of BFtS are discussed in Appendix A.

### v5

@@ 3 Introducing Bias via an Adversarial Missingness Process
The AMADB problem is NP-hard.
See the proof in the Appendix. We frame AMADB as an adversarial version of active learning where the goal is to strategically minimize the accuracy of the imputation by selecting observed sensitive attributes. More specifically, we apply a popular formulation of active learning as a coverage problem (Yehuda et al., 2022; Ren et al., 2021). Theorem 1 can be interpreted as a positive result, as it shows that, in theory, a simpler surrogate of the adversary’s objective is still hard to optimize.
[…]
$\Delta DP_{a}=|p(\hat{y}=1|a=0)$- $p(\hat{y}=1|a=1)|$
We want to design a heuristic where the goal of the adversary is to choose missing values to maximize the difference between the true demographic parity $\Delta DP_{s}$ and empirical demographic parity $\Delta DP_{\hat{s}}$:
$\max_{\hat{s}}~\Delta DP_{s}-\Delta DP_{\hat{s}}$
$=\max_{\hat{s}}~|p(\hat{y}=1|s=0)$- $p(\hat{y}=1|s=1)|-|p(\hat{y}=1|\hat{s}=0)-p(\hat{y}=1|\hat{s}=1)|$
We assume the minority class is underrepresented, i.e. $(\forall a,~|p(\hat{y}=1|a=0)\geq p(\hat{y}=1|a=1)|)$:
$\max_{\hat{s}}~p(\hat{y}=1|s=0)$- $p(\hat{y}=1|s=1)-p(\hat{y}=1|\hat{s}=0)+p(\hat{y}=1|\hat{s}=1)$
$=\max_{\hat{s}}~p(\hat{y}=1|s=0)$- $p(\hat{y}=1|\hat{s}=0)+p(\hat{y}=1|\hat{s}=1)-p(\hat{y}=1|s=1)$
$=\max_{\hat{s}}~p(\hat{y}=1|s=0)$- $p(\hat{y}=1|\hat{s}=0)+p(\hat{y}=0|s=1)-p(\hat{y}=0|\hat{s}=1)$
[…]
Definition 3.
The degree bias assumption: Given nodes $u,v\in\mathcal{V}$ where $deg(u)>deg(v)$, we assume that $p(s_{v}\neq\hat{s}_{v})>p(s_{u}\neq\hat{s}_{u})$.
The degree bias assumption has been supported by both theoretical and empirical results in the literature for both homophilic and non-homophilic graphs (Tang et al., 2020; Liu et al., 2023; Ju et al., 2024; Subramonian et al., 2024). Moreover, low-degree nodes are known to be more vulnerable to attacks than high-degree nodes (Zügner et al., 2018; Zügner & Günnemann, 2019). Therefore, our adversarial missingness process simply selects low-degree nodes to have missing sensitive values. We call the heuristic based on the degree bias assumption the ‘degree’ heuristic.
We also consider another heuristic to evaluate the effectiveness of the degree one. We select missing values uniformly at random from the set $\mathcal{V}_{miss}=\{v\in\mathcal{V}|(y_{v}=1\wedge s_{v}=0)\vee\ (y_{v}=0\wedge s_{v}=1)\}$. If the desired number of missing nodes exceeds $|\mathcal{V}_{miss}|$, we draw the remaining $|\mathcal{V}\setminus\mathcal{V}_{S}|-|\mathcal{V}_{miss}|$ samples uniformly at random from $\mathcal{V}\setminus\mathcal{V}_{miss}$. We call the resulting heuristic ‘targeted’.
To assess the effect of degree-based adversarial missingness, we compare imputation under adversarial degree, targeted and random missingness. Figure 2 reports correlations between sensitive attributes and class labels across datasets. Using a G Convolutional Network (GCN) (Kipf & Welling, 2017) for independent imputation, we observe that with less than 50% sensitive data, adversarial missingness consistently underestimates the true bias in the dataset, while random missingness and targeted missingness perform slightly better (Figures 2, top row). In contrast, BFtS (Figures 2, bottom row), introduced in the next section, rarely underestimates bias. These results highlight that off-the-shelf imputation can mislead fair graph learning by underestimating bias in the dataset, particularly under adversarial missingness. Among the three missingness processes, the degree heuristic exhibits the largest bias discrepancy between imputed and original values (i.e., is more adversarial). Therefore, we focus our analysis on the degree heuristic.
Figure 2: Empirical evaluation of three missingness processes and their effect on bias estimation. Degree-based adversarial missingness most effectively drives an independently trained GCN imputation model to underestimate bias, compared with random and targeted missingness. The final column in each matrix reports the true bias, where lower values indicate stronger underestimation. NBA shows the largest correlation gap (up to a 433% discrepancy when only 20% of protected attributes are observed) under the degree heuristic. We also include results for BFtS (Section 4), our 3-player adversarial imputation framework.
4 Fairness-Aware Adversarial Missing Data Imputation
@@
We introduce BFtS (Better Fair than Sorry), a 3-player adversarial framework for fair GNN training under adversarially missing sensitive data. The fair GNN is trained jointly with two adversaries–one to predict sensitive attributes based on GNN embeddings and another to impute missing values that minimize fairness, ensuring that fairness is evaluated against worst-case imputations.
We motivate the worst-case assumption using distributionally robust optimization (Ben-Tal et al., 2013; Mandal et al., 2020; Namkoong & Duchi, 2016; Shafieezadeh Abadeh et al., 2015). Let $\mathcal{P}_{s}$ be the sensitive attribute distribution. We can express the fairness objective in terms of the expected bias.
$\theta^{*}_{class}=\arg\min_{\theta_{class}}\mathcal{L}_{class}+\alpha\mathbb{E}_{{s}\sim\mathcal{P}_{s}}[\mathcal{L}_{bias}]$
However, under adversarial missingness, the true distribution of sensitive values cannot be accurately estimated from the observations (Mohan et al., 2013; Tian, 2017). To handle this uncertainty, let us define an uncertainty set $\mathcal{U}$ of plausible distributions for $s$. This leads to the worst-case imputation:
$\theta^{*}_{class}=\arg\min_{\theta_{class}}\mathcal{L}_{class}+\alpha\max_{u\in\mathcal{U}}\mathbb{E}_{{s}\sim u}[\mathcal{L}_{bias}]$
@@ 4.1 Proposed Model (BFtS)
Our goal is to learn a group fair and accurate GNN $\hat{y}=f_{class}(\mathcal{G},\mathcal{X})$ for node classification with adversarially missing sensitive values. The proposed solution is model-agnostic (e.g., GNN) and based on adversarial learning. It formulates the imputation model as a second adversary of the GNN. Figure 3 depicts the flow diagram of the proposed model. The model has three primary components: a missing sensitive attribute imputation GNN $f_{imp}$, a node classification GNN $f_{class}$, and a sensitive attribute Deep Neural Net (DNN) prediction $f_{bias}$. $f_{class}$ takes $\mathcal{X}$ and $\mathcal{G}$ as inputs and predicts the node labels. $f_{bias}$ is an adversarial neural network that attempts to estimate the sensitive information from the final layer representations of $f_{class}$ to assess the bias. More specifically, $f_{class}$ is biased if the adversary $f_{bias}$ can accurately predict the sensitive attribute information from the representations of $f_{class}$. The model $f_{imp}$ predicts the missing sensitive attributes by taking $\mathcal{X}$ and $\mathcal{G}$ as inputs and generating the missing sensitive attributes $\hat{si}$. The goal of $f_{imp}$ is to generate sensitive values that minimize the fairness of $f_{class}$, and, thus, it works as a second adversary to $f_{class}$.
Figure 3: 3-player framework for fair GNN training with missing data imputation (BFtS). $f_{class}$ generates node representations $\mathbf{h}_{v}$ by minimizing the classification loss $\mathcal{L}_{class}$ (Eqn. 1) and the maximizing sensitive attribute prediction loss $\mathcal{L}_{bias}$ (Eqn. 3). $f_{bias}$ predicts sensitive attributes using representations from $f_{class}$ by minimizing $\mathcal{L}_{bias}$. $f_{imp}$ predicts missing values by minimizing the imputation loss $\mathcal{L}_{imp}$ (Eqn. 2) and maximizing $\mathcal{L}_{bias}$. $\hat{y}$, $\hat{si}$, $\hat{sa}$ are predictions from $f_{class}$, $f_{imp}$ and $f_{bias}$, respectively.
@@ 4.1.1 Player architectures
GNN classifier $f_{class}$:
Node classification model $\hat{y}_{v}=f_{class}(x_{v},\mathcal{G})$ implemented using a GNN. We assume that $f_{class}$ does not apply sensitive attributes but uses other attributes in $\mathcal{X}$ that might be correlated with sensitive ones. The goal of $f_{class}$ is to achieve both accuracy and fairness. To improve fairness, $f_{class}$ tries to minimize the loss of adversary $f_{bias}$.
Sensitive attribute predictor (Adversary 1) $f_{bias}$:
Neural network that uses representations $\mathbf{h}_{v}$ from $f_{class}$ to predict sensitive attributes as $\hat{sa}_{v}=f_{bias}(\mathbf{h}_{v})$. $f_{class}$ is fair if $f_{bias}$ performs poorly.
Missing data imputation GNN (Adversary 2) $f_{imp}$:
Predicts missing sensitive attributes as $\hat{si}_{v}=f_{imp}(x_{v},\mathcal{G})$. However, besides being accurate, $f_{imp}$ plays the role of an adversary to $f_{class}$ by predicting values that maximize the accuracy of $f_{bias}$.
@@ 4.1.2 Loss Functions
Node Classification:
We apply the cross-entropy loss to learn $f_{class}$ as follows:
$\mathcal{L}_{class}=-\frac{1}{|{\mathcal{V}_{L}}|}\sum_{v\in{\mathcal{V}_{L}}}\hat{y}_{v}log(\hat{y}_{v})+(1-\hat{y}_{v})log(1-\hat{y}_{v}).$ (1)
Sensitive Attribute Imputation:
Because sensitive attributes tend to be imbalanced, $f_{imp}$ applies the Label-Distribution-Aware Margin (LDAM) loss Cao et al. (2019). Let $\hat{si}=f_{imp}(X,G)$ be the one-hot encoded predictions and $\hat{si}^{s}_{v}$ be the prediction for $s_{v}$. The LDAM loss of $f_{imp}$ is defined as:
$\mathcal{L}_{imp}=\frac{1}{|{\mathcal{V}_{S}}|}\sum_{v=1}^{|\mathcal{V}_{S}|}-log\frac{e^{\hat{si}^{s}_{v}-\Delta^{s}}}{e^{\hat{si}^{s}_{v}-\Delta^{s}}+\sum_{k\neq s}e^{\hat{si}^{k}_{v}}}.$ (2)
@@
The LDAM loss is a weighted version of the negative log-likelihood loss. Intuitively, since $\Delta^{j}$ is larger for smaller values of $n_{j}$, and thus it ensures a higher margin for the smaller classes.
Sensitive attribute imputation with no sensitive information: BFtS can generate fair outputs when very little or no sensitive information is provided by letting $f_{imp}$ impute the sensitive values from the ground truth training labels $y$ using the LDAM loss Du et al. (2021). More specifically, we replace $\mathcal{V}_{S}$ with $\mathcal{V}_{L}$ and $s$ with $y$ for each node $v\in\mathcal{V}_{L}$ in Eqn 2. The reasoning is that the worst-case fairness model generally assigns a more desired outcome to the non-sensitive group and a less desired outcome to the sensitive one. Therefore, the nodes predicted as the minority class will fall into the sensitive group, and vice versa. This is consistent with the worst-case assumption of BFtS.
Sensitive Attribute Prediction:
Given the sensitive information for $\color[rgb]{1,0,0}\mathcal{V}_{S}$}, we first replace $\hat{si}_{v}$ by ${s}_{v}$ for $v\in{\mathcal{V}_{S}}$ and thereby generate $\hat{s}$. The parameters of $f_{bias}$ are learned using:
$\mathcal{L}_{bias}=\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{bias}(\mathbf{h})]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{bias}(\mathbf{h}))]$ (3)
For simplicity, we consider a binary sensitive attribute in this paper. However, BFtS extends to non-binary or continuous sensitive attributes by appropriately generalizing the imputation loss $\mathcal{L}_{\text{imp}}$ and the bias loss $\mathcal{L}_{\text{bias}}$ to multi-class or continuous settings.
@@ 4.1.3 Learning the parameters of BFtS
Let $\theta_{class},\theta_{bias}$, and $\theta_{imp}$ be the parameters of $f_{class}$, $f_{bias}$, and $f_{imp}$, respectively, which are learned via a 3-player adversarial scheme described in Figure 3. Parameters $\theta_{class}$ are optimized as:
$\theta_{class}^{*}=\arg\min_{\theta_{class}}\mathcal{L}_{class}+\alpha\mathcal{L}_{bias}.$ (4)
@@
The parameters of the sensitive attribute predictor $\theta_{bias}$ are learned by maximizing $\mathcal{L}_{bias}$:
$\theta_{bias}^{*}=\arg\max_{\theta_{bias}}\mathcal{L}_{bias}.$ (5)
To learn the parameters $\theta_{imp}$ of the imputation model to generate predictions that are accurate and represent the worst-case scenario for fairness, we apply the following:
$\theta_{imp}^{*}=\arg\min_{\theta_{imp}}\mathcal{L}_{imp}-\beta\mathcal{L}_{bias}.$ (6)
@@
$\begin{split}\min_{\theta_{class}}\max_{\theta_{imp},\theta_{bias}}\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=1)}[\log f_{bias}(\mathbf{h})]+\mathbb{E}_{\mathbf{h}\sim p(\mathbf{h}|\hat{s}=0)}[\log(1-f_{bias}(\mathbf{h}))].\end{split}$ (7)
The training algorithm and time complexity analysis of BFtS are discussed in Appendix.